OpenStudy (anonymous):

where do i type exponent 2 on the keyboard

OpenStudy (anonymous):

^2

OpenStudy (lgbasallote):

use ^2

OpenStudy (anonymous):

do i enter it?

OpenStudy (anonymous):

shift-6

OpenStudy (lgbasallote):

it's just how people denote it for example x squared is x^2

OpenStudy (anonymous):

shift 6 type 2

OpenStudy (anonymous):

^6

OpenStudy (anonymous):

like taht?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

it doesnt look like an exponent though

OpenStudy (anonymous):

that exponent sign

OpenStudy (anonymous):

*thats

OpenStudy (anonymous):

how to solve this (-2)^2-4(-2)-4

OpenStudy (lgbasallote):

\(x^2\)

OpenStudy (anonymous):

how u do that

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

so confusing

OpenStudy (lgbasallote):

\[\text{\[x^2\]}\] type exacty what i typed

OpenStudy (anonymous):

cant find that first dash or slash

OpenStudy (anonymous):

nvm found it

OpenStudy (anonymous):

cant you do x^{2}

OpenStudy (anonymous):

or sumthing like that ^^

OpenStudy (anonymous):

\[x^2\]

OpenStudy (anonymous):

yaya

OpenStudy (anonymous):

how do i do just exponent 2 without the x?

OpenStudy (lgbasallote):

what do you mean

OpenStudy (anonymous):

(-2)^2-4(-2)-4 = 4-4(-2)-4=4+8-4= 8 ... so 8 is your FINAL answer!!!!

OpenStudy (anonymous):

8

OpenStudy (richyw):

In basically every math program and latex you use base^{exponent} or if the exponent is a single number just base^exponent. In microsoft word just do control shift plus, or control plus for superscripts and subscripts (can't remember which is which)

OpenStudy (richyw):

\[^2\]is just the exponent without the base, although it makes no sense and all you get is a smaller 2. Just type ^2